OpenStudy (anonymous):

Solve the compound inequality 6b < 24 or 4b + 12 > 4. b < 4 or b > −2 b < 4 or b > 4 b < 4 or b > 2 b > 4 or b < −2

OpenStudy (mathstudent55):

Solve each inequality and keep the word "or" in between them. \(6b < 24\) or \(4b + 12 > 4\) \(\dfrac{6b}{6} \lt \dfrac{24}{6} \) or \(4b + 12 -12 \gt 4 - 12\) \(b \lt 4\) or \(4b \gt -8 \) \(b \lt 4\) or \(\dfrac{4b}{4} \gt \dfrac{-8}{4} \) \(b \lt 4\) or \(b \gt -2\) Now look at the two inequalities. If b < 4 or b > -2, that means b can be any number. The solutions is all real numbers.

OpenStudy (anonymous):

why dont you flip the sign

OpenStudy (mathstudent55):

The sign in an inequality only flips direction when you multiply both sides or divide both sides by a negative number. In this case, we never did that.

OpenStudy (mathstudent55):

For example: \(-5x \le 30 \) Here you need to divide both sides by -5, obviously a negative number, so you get this: \(\dfrac{-5x}{-5} \ge \dfrac{30}{-5} \) \(\dfrac{\cancel{-5}x}{\cancel{-5}1} \ge -6 \) \(x \ge -6 \)

OpenStudy (anonymous):

the answer is A (b<4 or b>-2)

OpenStudy (mathstudent55):

Yes, it's A.